# Practical approaches for crystal structure predictions with inpainting generation and universal interatomic potentials

### v1

@@ III.4 Example: Li-Si alloys
To demonstrate the effectiveness of identifying low-energy polymorphs, we performed structure generation for Li-rich phases (LixSiy, where $x>y$, a composition range known to contain many stable phases [51]) following the workflow outlined in Fig. 3 (see Methods). After CHGGen generation and r2SCAN-DFT calculations, we constructed the phase diagram using DFT formation energies from both MP structures and the generated structures. In Fig. 5a, green circles represent new stable polymorphs identified by CHGGen, while blue squares indicate the reported stable MP structures. We calculated decomposition energies using the MP phase diagram (without the on-the-hull structure of Li5Si2). Negative values of $E_{d}$ therefore indicate compounds that break the existing MP convex hull. Notably, CHGGen successfully predicted Li5Si2 (R-3m, $E_{d}=-7$ meV/atom) as a thermodynamically stable polymorph beyond the known MP stable structures (Fig. 5b). Interestingly, this structure had also been identified through previous studies, including work by Tipton et al. [52] using genetic algorithms and Morris et al. [53] using random structure search coupled with DFT calculations.
Using generative models, one can also identify metastable polymorphs with low decomposition energies in related compositions. For example, structures for compositions corresponding to Li5Si2 ($E_{d}=-0.006$ eV/atom), Li2Si ($E_{d}=0.003$ eV/atom), and Li4Si ($E_{d}=0.010$ eV/atom) with C2/m spacegroup are shown in Fig. 5c. These low-energy polymorphs with different local structure motifs provide a more detailed mapping of the potential energy landscape, which benefits the training of MLIPs to understand the Si network aggregation and its effect on Li transport kinetics [54]. The case study illustrates how practitioners can benefit from combining diffusion models and uMLIPs to explore continuous unknown chemical spaces, such as those encountered in alloy design.
@@ IV Discussion
Based on this concept, diffusion models with GNNs represent a logical framework for generating reasonable local structure motifs (e.g., atomic bonding patterns), which proves valuable when optimizing local atomic arrangements in CSP problems. Nonetheless, their inherent locality bias limits their ability to capture long-range periodic orders. Gong et al. [46] revealed that state-of-the-art GNNs fall short of accurately capturing the periodicity of crystal structures, i.e., lattice parameters. This fundamental limitation explains the diminished performance of GNN-based diffusion models when generating structures with large unit cells, where long-period crystallinity is significant but hard to capture.
Host-structure-guided generation with inpainting presents another practical approach to addressing symmetric structure generation. To evaluate the practical advantage of inpainting, we compared the success rates of obtaining symmetric crystals across different approaches in Zn-P-S and Li-Si chemical spaces: MatterGen (green), CHGGen with unconditional generation (blue), and CHGGen with inpainting generation (orange). MatterGen demonstrates superior performance compared to our baseline model for Li-Si alloys (Fig. 6b), highlighting the importance of explicitly modeling lattice diffusion in conjunction with atomic arrangements. However, for more complex systems such as Zn-P-S that contain polyanions, MatterGen performs less effectively when the number of atoms exceeds 10, with success rates lower than CHGGen with inpainting generation. Notably, MatterGen exhibits declining success rates as system size increases in both cases, suggesting that it is likely all GNN-based diffusion models face scalability challenges. This scalability constraint potentially limits their ability to predict complex structural prototypes (e.g., NASICON-type frameworks) across various solid-state materials.
@@
Although CHGGen has not yet achieved industry-level performance as MatterGen in stability or general symmetry success rate, it shows clear improvements in generating symmetric crystal structures compared to unconditional generation methods. Assessing crystallographic symmetry in generated structures is critical for CSP, as discovering new structural prototypes is a key step toward materials discovery. These prototypes can subsequently guide elemental substitution strategies [15] or evolutionary searches [61] for compositional optimization. The inpainting-based generation approach samples from the conditional distribution of unknown structural components within a given framework, yielding more well-defined local atomic arrangements than de novo sampling from a fully unconstrained distribution. Importantly, this inpainting strategy is simple to implement, as both conditional and unconditional generation operate within the same unified model—distinguished only by the masking strategy used during the reverse diffusion process. This modular design enables seamless integration with existing foundational generative models and provides a flexible mechanism for enforcing symmetry constraints during structure generation.
@@
Finally, while our framework demonstrates useful augmentation to diffusion-based generative models, the current symmetry refinement approach remains preliminary as it relies on spglib by simply increasing the tolerance threshold. As a proof-of-concept, this method predominantly yields structures with moderate symmetry (e.g., C2, Cm in monoclinic systems), which limits the discovery of novel structure prototypes. Looking forward, several promising approaches have emerged for novel framework generation, including symmetry-constrained diffusion [23] and prototype-based generation using Wyckoff position-based representations [65, 66, 67]. The integration of these advanced symmetry handling strategies with CHGGen could enhance the discovery of crystal structures particularly with intercalation chemistry.
In summary, we present CHGGen as an integrated framework that combines unconditional and inpainting generation with uMLIP optimization for CSP. While challenges in scaling complexity persist, the inpainting method provides a useful approach for generating symmetric crystals and incorporating intercalants into existing database structures. We anticipate broad adoption of this framework as the modular design of the inpainting methodology enables seamless integration with emerging diffusion models, which will ultimately accelerate materials discovery across diverse chemical spaces.
@@ Methods
Machine learning interatomic potential. For the generated structures, we used the pre-trained CHGNet as an efficient and accurate calculator to evaluate the energy and interatomic forces. The pre-trained CHGNet achieves mean-absolute errors of 30 meV/atom for energy and 70 meV/Å for interatomic forces on the MPtrj dataset against the DFT at GGA/GGA$+U$ level of accuracy [12]. The structure relaxations were optimized by the FIRE optimizer [73] over the potential energy surface provided by CHGNet, where the atom positions, cell shape, and cell volume were optimized to reach converged interatomic forces of 0.1 eV/Å[12].

### v2

@@ III.4 Example: Li-Si alloys
To demonstrate the effectiveness of identifying low-energy polymorphs, we performed structure generation for Li-rich phases (LixSiy, where $x>y$, a composition range known to contain many stable phases [52]) following the workflow outlined in Fig. 3 (see Methods in SI). After CHGGen generation and r2SCAN-DFT calculations, we constructed the phase diagram using DFT formation energies from both MP structures and the generated structures. In Fig. 5a, green circles represent new stable polymorphs identified by CHGGen, while blue squares indicate the reported stable MP structures. We calculated decomposition energies using the MP phase diagram (without the on-the-hull structure of Li5Si2). Negative values of $E_{d}$ therefore indicate compounds that break the existing MP convex hull. Notably, CHGGen successfully predicted Li5Si2 (R-3m, $E_{d}=-7$ meV/atom) as a thermodynamically stable polymorph beyond the known MP stable structures (Fig. 5b). Interestingly, this structure had also been identified through previous studies, including work by Tipton et al. [53] using genetic algorithms and Morris et al. [54] using random structure search coupled with DFT calculations.
Using generative models, one can also identify metastable polymorphs with low decomposition energies in related compositions. For example, structures for compositions corresponding to Li5Si2 ($E_{d}=-0.006$ eV/atom), Li2Si ($E_{d}=0.003$ eV/atom), and Li4Si ($E_{d}=0.010$ eV/atom) with C2/m spacegroup are shown in Fig. 5c. These low-energy polymorphs with different local structure motifs provide a more detailed mapping of the potential energy landscape, which benefits the training of MLIPs to understand the Si network aggregation and its effect on Li transport kinetics [55]. The case study illustrates how practitioners can benefit from combining diffusion models and foundation potentials to explore continuous unknown chemical spaces, such as those encountered in alloy design.
@@ IV Discussion
Based on this concept, diffusion models with GNNs represent a logical framework for generating reasonable local structure motifs (e.g., atomic bonding patterns), which proves valuable when optimizing local atomic arrangements in CSP problems. Nonetheless, their inherent locality bias limits their ability to capture long-range periodic orders. Gong et al. [47] revealed that state-of-the-art GNNs fall short of accurately capturing the periodicity of crystal structures, i.e., lattice parameters. This fundamental limitation explains the diminished performance of GNN-based diffusion models when generating structures with large unit cells, where long-period crystallinity is significant but hard to capture. This locality bias is exacerbated for species that can adapt to diverse local environments (e.g., Li), which have a broad distribution of stable coordination geometries. The unconditional generation defaults to producing a disordered ”mosaic” of these competing local motifs rather than a coherent crystal structure.
The practical advantage of host-structure-guided generation with inpainting is to augment the symmetric structure generation. To evaluate this, we compared the success rates of obtaining symmetric crystals across different approaches in Zn-P-S and Li-Si chemical spaces: MatterGen (green), CHGGen with unconditional generation (blue), and CHGGen with inpainting generation (orange). MatterGen demonstrates superior performance compared to our baseline model for Li-Si alloys (Fig. 6b), highlighting the importance of explicitly modeling lattice diffusion in conjunction with atomic arrangements. However, for more complex systems such as Zn-P-S that contain polyanions, MatterGen performs less effectively when the number of atoms exceeds 10, with success rates lower than CHGGen with inpainting generation. This highlights the importance of inpainting generation when dealing with chemical systems involving polyanions. Notably, MatterGen exhibits declining success rates as system size increases in both cases, suggesting that it is likely all GNN-based diffusion models face scalability challenges. This scalability constraint potentially limits their ability to predict complex structural prototypes (e.g., NASICON-type frameworks) across various solid-state materials.
@@
Although CHGGen has not yet achieved industry-level performance as MatterGen in stability or general symmetry success rate, it shows clear improvements in generating symmetric crystal structures compared to unconditional generation methods. Assessing crystallographic symmetry in generated structures is critical for CSP, as discovering new structural prototypes is a key step toward materials discovery [23]. These prototypes can subsequently guide elemental substitution strategies [15] or evolutionary searches [62] for compositional optimization. The inpainting-based generation approach samples from the conditional distribution of unknown structural components within a given framework, yielding more well-defined local atomic arrangements than de novo sampling from a fully unconstrained distribution. Importantly, this inpainting strategy is simple to implement, as both conditional and unconditional generation operate within the same unified model—distinguished only by the masking strategy used during the reverse diffusion process. This modular design enables seamless integration with existing foundational generative models and provides a flexible mechanism for enforcing symmetry constraints during structure generation.
@@
Finally, while our framework demonstrates useful augmentation to diffusion-based generative models, the current symmetry refinement approach remains preliminary as it relies on spglib by simply increasing the tolerance threshold. As a proof-of-concept, this method predominantly yields structures with moderate symmetry (e.g., C2, Cm in monoclinic systems), which may limit the discovery of novel structure prototypes (see Fig. S1 and S2). Looking forward, several promising approaches have emerged for novel framework generation, including symmetry-constrained diffusion [24] and prototype-based generation using Wyckoff position-based representations [66, 67, 68]. The integration of these advanced symmetry handling strategies with CHGGen could enhance the discovery of crystal structures particularly with intercalation chemistry.
In summary, we present CHGGen as an integrated framework that combines unconditional and inpainting generation with foundation potential optimization for CSP. While challenges in scaling complexity persist, the inpainting method provides a useful approach for generating symmetric crystals and incorporating intercalants into existing database structures. We anticipate broad adoption of this framework as the modular design of the inpainting methodology enables seamless integration with emerging diffusion models, which will ultimately accelerate materials discovery across diverse chemical spaces.
@@ Methods
Machine learning interatomic potential. For the generated structures, we used the pre-trained CHGNet as an efficient and accurate calculator to evaluate the energy and interatomic forces. The pre-trained CHGNet achieves mean-absolute errors of 30 meV/atom for energy and 70 meV/Å for interatomic forces on the MPtrj dataset against the DFT at GGA/GGA+U+U level of accuracy [12]. The structure relaxations were optimized by the FIRE optimizer [74] over the potential energy surface provided by CHGNet, where the atom positions, cell shape, and cell volume were optimized to reach converged interatomic forces of 0.1 eV/Å[12].